OpenStudy (anonymous):

Convert the function into vertex form. Show your work. y = 2(x - 3)(x + 4)

OpenStudy (anonymous):

y= 2(x^2-3x+4x -12) y = 2(x^2+x-12) ax^2 + 2abx +b^2 2ab = 1 a = 1 so b = 0.5 so b^2 = 0.25 y = 2(x^2 +x +0.25 -12.25) y = 2(x+0.5)^2 - 24.5